OpenStudy (anonymous):

Prove x2+y2=25 equation of a circle?

OpenStudy (anonymous):

it is in the form of (x-h)^2 + (y-k)^2 = r^2

OpenStudy (anonymous):

or just draw it :D

OpenStudy (precal):

you could also, state that the center is (0,0) and the radius is 5

OpenStudy (anonymous):

Solve √[(x - 0)^2 + (y - 0)^2] = 5. This comes from the definition of a circle. The center is (0, 0) and the radius is 5, so the distance from any point on the circle (x,y) to the center is 5.

OpenStudy (anonymous):

@zaibali.qasmi Yours thoughts are so similar :P https://uk.answers.yahoo.com/question/index?qid=20100207094206AAmD62U

OpenStudy (precal):

but why take the square root? this is already in standard form of a circle

OpenStudy (anonymous):

yupp,,,i study it there an think it is simplest form to understand the asker..thats why..!!

OpenStudy (precal):

what level of math is this? If it is geometry or algebra on a high school level, I would think stating that this is the standard form of a circle would suffice. If this is college level, then you more likely have to use proofs. Just wondering....

myininaya (myininaya):

It does say prove.

OpenStudy (precal):

true but it is also, important to find out what level it is? sometimes justify, explain and prove are misuse on the high school or secondary level.

OpenStudy (jdoe0001):

I see everyone's posting BUT keepsecure93 's, wonder why

OpenStudy (precal):

because keepsecure93 is not online (perhaps)

OpenStudy (jdoe0001):

well, his nick is

OpenStudy (jdoe0001):

anyhow -> http://www.mathwarehouse.com/geometry/circle/images/equation-of-circle/example-2-equation-of-circle.png <--- surely you've covered this by now

OpenStudy (anonymous):

DEAR MEMBERS.I AM A STUDENT OF 8THCLAS AND DIDNOT KNO MUCH ENGLISH .I COPY QUESTION FROM BOK

OpenStudy (jdoe0001):

hmm that still doesn't exempt you from covering the material on circles or replying for that matter

OpenStudy (precal):

Is the book a geometry or algebra?

OpenStudy (precal):

geometry would want a proof algebra probably just want you to recognize that this is standard form of a conic and that you can say, yes it is a circle with the center (0,0) radius 5

OpenStudy (anonymous):

BASE GEOMETRY

OpenStudy (jdoe0001):

likely just a match-up against the standard circle formula

OpenStudy (precal):

well then I stand by my response

OpenStudy (anonymous):

MNE 2SRA SWAL KESY POCHN